(3 4 -1 1

1 _1 3 1

4 _3 11 2. Reduce to row reduced echelon form


1
Expert's answer
2021-01-21T04:11:34-0500

"\begin{pmatrix}
 3 & 4 & -1 & 1\\
 1 & -1 & 3& 1\\
 4 & -3 & 11 & 2
\end{pmatrix}"


Let's divide the 1st row by 3


"\begin{pmatrix}
 1 & 4/3 & -1/3 & 1/3\\
 1 & -1 & 3& 1\\
 4 & -3 & 11 & 2
\end{pmatrix}"


From the 2nd row we have to subtract the first row; from the 3rd row we have to subtract the 4 times first row.


"\begin{pmatrix}
 1 & 4/3 & -1/3 & 1/3\\
 0 & -7/3 & 10/3& 2/3\\
 0 & -25/3 & 37/3 & 2/3
\end{pmatrix}"


Let's divide the second row by -7/3


"\begin{pmatrix}
 1 & 4/3 & -1/3 & 1/3\\
 0 & 1 & -10/7& -2/7\\
 0 & -25/3 & 37/3 & 2/3
\end{pmatrix}"


To the third row we have to add the second row multiplied by 25/3


"\begin{pmatrix}
 1 & 4/3 & -1/3 & 1/3\\
 0 & 1 & -10/7& -2/7\\
 0 & 0 & 3/7 & -12/7
\end{pmatrix}"


Let's divide the third row by 3/7


So the answer is: "\begin{pmatrix}
 1 & 4/3 & -1/3 & 1/3\\
 0 & 1 & -10/7& -2/7\\
 0 & 0 & 1 & -4
\end{pmatrix}"